OpenStudy (anonymous):

x+a=b/3x and b does not equal 3 then x=

OpenStudy (anonymous):

\[x+a=\frac{b}{3x}\] \[3x^2+3ax=b\] \[3x^2+3ax-b=0\] then i guess the quadratic formula

OpenStudy (eyust707):

do you know the quadratic formula?